OpenStudy (123ab456c):

The area of a triangle is 56.4 cm2 and the base is 12 cm. What is the height of the triangle?

OpenStudy (stormer15):

\(56.4=\dfrac{12*b}{2}\) first step, multiply 56.4 * 2 = ?

OpenStudy (123ab456c):

is it 4.7

OpenStudy (stormer15):

the final answer?

OpenStudy (saritamust):

9.4 cm?

OpenStudy (stormer15):

ye

OpenStudy (michele_laino):

hint: we have to do this computation: \[\Large height = \frac{{2 \times 56.4}}{{12}} = ...?\]